You have a 10 ohm resistor connected in series with an ammeter. The voltage applied to the whole circuit is 1.2 volts. At the same time you have a voltmeter connected across the 10 ohm resistor and voltmeter reads 0.9 volts. The circuit draws a total of 0.5 amps. Can you determine resistance of the ammeter?
Answer #1

The potential drop across the voltemeter is 0.9V the memaining voltage across ammeter is 1.2-0.9= 0.3V

this is the potential drop across the ammeter. the resistance of ammeter is r=V/I = 0.3/0.5 = 0.6ohms

voltmeter has high internnal resistance where as the ammeter has very low internal resistance.